Show that (tanx + cotx)cos2x = 2cot2x , hence solve the equation (tanx + cotx)cos2x = 2 , for 0
Answer by Alan3354(69443)

Show that (tanx + cotx)cos2x = 2cot2x = 2(cos(2x)/sin(2x))
(tanx + cotx) = 2/sin(2x)
sinx/cosx + cosx/sinx = 2/2sinxcosx = 1/sinxcosx
Multiply by sinx*cosx
sin^2x + cos^2x = 1 An identity
--------------------------
solve the equation (tanx + cotx)cos2x = 2 , for 0
what 0